Cost function=400Q-50Q^2+5Q^3

What is the level of output at which the production of the uniform minimize the company's ATC


1
Expert's answer
2021-05-13T17:43:17-0400

"TC = 400Q-500Q^2+5Q^3"

"ATC =\frac{TC}{Q}=400-500Q+5Q^2"

"\text{find min ATC:}"

"ATC'=(400-500Q+5Q^2)'=-500+10Q"

"ATC'=0;-500+10Q=0;Q=50"

"ATC'(49)=-500+490=-10<0"

"ATC'(51)=-500+510=10>0"

"\text{hence } Q=50 \text{minimum point of function }ATC"

"\text{Answer: }50"